OpenStudy (anonymous):

1- Alicia and Jamal completed the following proof: Given: 3x+7=x-5 Proof: x=-6 Alicia's Method: Statements- Reasons- 3x+7=x-5 Given 2x+7=-5 Subtraction Property of Equality 2x=-12 Subtraction Property of Equality x=-6 Division Property of Equality Jamal's Method: Statements- Reasons- 3x+7=x-5 Given 7=-2x-5 Subtraction property of equality 12=-2x Addition property of equality

OpenStudy (danjs):

ok

OpenStudy (anonymous):

-6=x Division property of equality Who completed the proof correctly? A- Alicia only B- Jamal Only C- Both Alicia and Jamal D- neither Alicia nor Jamal 2- which of the following statements are true? Multiple Choice A- If two angles form a linear pair, them the angles are supplementary. B- If two angles are right angles, then the angles are complementary. C- If two angles have the same measure, then the angles are congruent. D- If two angles are supplementary, then the angles are acute. 3-Use the given property to complete the statement. Distributive property 3(x-1)=3x-_____? A- -1 B - -3 C- 1 D- 3 4- Statements: Reasons: 1/2x-5=10 Given 2(1/2x-5)=20 ? x-10=20 ? x=30 ? Part A: What is the first missing reason? A- Distributive property B- Multiplication property of equality C- Addition property of equality D- Substituition property of equality Part B: What is the second missing reason? A Distributive Property B Multiplication Property of Equality C Addition Property of Equality D Substitution Property of Equality Part C: What is the third missing reason? A Distributive Property B Multiplication Property of equatlity C Addition property of equality D Substitution property of equality Thanks for helping me.

OpenStudy (danjs):

haha, yuou paste a book

OpenStudy (anonymous):

What I think Q1- C Q2- C and D Q3- A Q4- A-c B- A B- B

OpenStudy (anonymous):

I mean C is B

OpenStudy (danjs):

for the first proof prob. ; both look right. There really only is an addition property and an multiplication property i think... subtraction is just addition of negative numbers division is just multiplication by fractions..

OpenStudy (anonymous):

thanks

OpenStudy (danjs):

2) You say C and D - Two angles of the same measure are congruent, yes true -- Not so, Supplimentary angles total to 180 degrees, If the angles are acute (less than 90), then two angles are NOT supplimentary, need one acute and one obtuse, or two right angles

OpenStudy (danjs):

**2a. Linear pair of angles are supplimentary...yes, linear pair of angles span a straight line angle, or 180 deg, just the angle you need for them to be considered supplimentary

OpenStudy (danjs):

for 2b, false , two right angles add to 180 deg, complimentary angles total 90 deg

OpenStudy (anonymous):

so for number 2 it would be B and D?